A flywheel of mass 400 kg and radius of gyration 1 m losses its speed from 300 r.p.m. to 240 r.p.m. m 120 seconds. Determine the retardmg torque acting on it.


(Ans, 20.8 N-m).


1
Expert's answer
2021-02-19T10:31:21-0500

Let's first find the angular acceleration of the flywheel:


α=ωfωit,\alpha=\dfrac{\omega_f-\omega_i}{t},α=240 revmin1 min60 s2π rad300 revmin1 min60 s2π rad120 s,\alpha=\dfrac{240\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ s}\cdot2\pi\ rad-300\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ s}\cdot2\pi\ rad}{120\ s},α=0.052 rads2.\alpha=-0.052\ \dfrac{rad}{s^2}.

The sign minus means that the flywheel is decelerates.

Next, let's calculate the moment of inertia of the flywheel:


I=mk2=400 kg(1 m)2=400 kgm2.I=mk^2=400\ kg\cdot(1\ m)^2=400\ kg\cdot m^2.

Finally, we can find the retarding torque acting on the flywheel from the formula:


τ=Iα=400 kgm2(0.052 rads2)=20.8 Nm.\tau=I\alpha=400\ kg\cdot m^2\cdot(-0.052\ \dfrac{rad}{s^2})=-20.8\ N\cdot m.

The magnitude of the retarding torque is 20.8 Nm20.8\ N\cdot m. The sign minus means that the torque acts in the opposite direction to the motion of the flywheel.